# Revisiting Applicable and Comprehensive Knowledge Tracing in Large-Scale Data

### v1

@@ Abstract
Knowledge Tracing (KT) is a fundamental component of Intelligent Tutoring Systems (ITS), enabling the modeling of students’ knowledge states to predict future performance. The introduction of Deep Knowledge Tracing (DKT), the first deep learning-based KT (DLKT) model, has brought significant advantages in terms of applicability and comprehensiveness. However, recent DLKT models, such as Attentive Knowledge Tracing (AKT), have often prioritized predictive performance at the expense of these benefits. While deep sequential models like DKT have shown potential, they face challenges related to parallel computing, storage decision modification, and limited storage capacity. To address these limitations, we propose DKT2, a novel KT model that leverages the recently developed xLSTM architecture. DKT2 enhances input representation using the Rasch model and incorporates Item Response Theory (IRT) for interpretability, allowing for the decomposition of learned knowledge into familiar and unfamiliar knowledge. By integrating this knowledge with predicted questions, DKT2 generates comprehensive knowledge states. Extensive experiments conducted across three large-scale datasets demonstrate that DKT2 consistently outperforms 17 baseline models in various prediction tasks, underscoring its potential for real-world educational applications. This work bridges the gap between theoretical advancements and practical implementation in KT. Our code and datasets will be available at https://github.com/codebase-2025/DKT2.
Revisiting Applicable and Comprehensive Knowledge Tracing in Large-Scale Data
Yiyun Zhou, Wenkang Han, Jingyuan Chenthanks: Corresponding author Zhejiang University {yiyunzhou, wenkangh, jingyuanchen}@zju.edu.cn
@@ 1 Introduction
The rapid expansion of educational data within Intelligent Tutoring Systems (ITS) Luckin and Holmes (2016) (e.g., AutoTutor Nye et al. (2014)) has exposed significant limitations in traditional machine learning approaches Bengio et al. (2013). In contrast, the advent of deep learning has introduced novel opportunities for addressing these challenges LeCun et al. (2015). A critical component of ITS is Knowledge Tracing (KT), which models students’ knowledge states and predicts future performance by analyzing their interaction data. Deep learning, with its advanced feature learning paradigm, offers enhanced modeling power and predictive accuracy in this context.
@@
Deep Knowledge Tracing (DKT) Piech et al. (2015) represents the first significant application of deep learning to KT, employing Long Short-Term Memory (LSTM) networks Hochreiter and Schmidhuber (1997) to capture the complexity of students’ learning processes. As a pioneering deep learning-based KT (DLKT) model, DKT has demonstrated superior predictive performance compared to traditional machine learning-based KT models (e.g., Bayesian Knowledge Tracing (BKT) Corbett and Anderson (1994)), offering notable advantages in applicability and comprehensiveness.
@@
DKT encodes students’ historical interactions to generate a comprehensive representation of their knowledge states (i.e., proficiency scores [^1] Proficiency scores range from 0 to 1, with higher values indicating greater knowledge and skill level. for each concept at each time step) and predicts future performance. However, recent DLKT models, such as the Attentive Knowledge Tracing (AKT) Ghosh et al. (2020), while excelling in predictive accuracy Liu et al. (2023b); Im et al. (2023); Huang et al. (2023); Yin et al. (2023), present limitations in applicability and comprehensiveness. Specifically, AKT requires both historical and future interactions as input (see Appendix  A.6.1 for input details), complicating its practical application since future responses are typically unavailable. Additionally, unlike DKT, AKT directly predicts scores on future questions without generating a comprehensive knowledge state, potentially weakening the correlations between different concepts and narrowing the definition of knowledge states in KT. Our review of 30 KT-related papers published in top AI/ML conferences and journals over the past decade (see Appendix A.1) reveals a trend where evaluation performance has been prioritized at the expense of practical applicability, risking a disconnect between theoretical advancements and real-world implementation.
@@
Deep sequential models like DKT have intrinsic limitations that may prevent them from achieving optimal performance. LSTM networks, for instance, face challenges in dynamically updating stored information and exhibit limited storage capacity due to their scalar cell state design. Moreover, their inherent sequential processing nature hinders parallelization, limiting their scalability to large datasets. The recently proposed xLSTM Beck et al. (2024), however, addresses these challenges by introducing two new variants: sLSTM, which improves LSTM’s storage decision by incorporating an exponential activation function, and mLSTM, which replaces scalar cell states with matrix memory for increased storage capacity and improved retrieval efficiency, while achieving full parallelization by abandoning memory mixing. Building on the strengths of xLSTM, we introduce DKT2, an enhanced DLKT model designed for greater applicability and comprehensiveness. DKT2 integrates the Rasch model Rasch (1993) from educational psychology to process historical interactions, using xLSTM for knowledge learning. The model then incorporates Item Response Theory (IRT) Lord (1952); Yen and Fitzpatrick (2006) to interpret the learned knowledge, differentiating between familiar and unfamiliar knowledge, and ultimately integrates this knowledge with predicted questions to generate comprehensive knowledge states.
@@
  • We introduce DKT2, a model built on xLSTM, adhering to rigorous input and output settings, and incorporating both the Rasch model for input and an interpretable IRT-based output module.
  • We conduct extensive experiments, including one-step prediction, multi-step prediction, and predictions with varying history lengths, across three large-scale datasets. Our findings demonstrate that DKT2 consistently outperforms 17 baseline models, with additional analysis on the impact of input settings and multi-concept predictions on KT performance.
@@ 2 Related Work
Since DKT Piech et al. (2015) first applied deep learning methods to the KT task nearly a decade ago, deep learning techniques have flourished in KT. Existing DLKT models can be categorized into the following 8 types:
  • Deep sequential models use recurrent structures to encode students’ chronologically ordered interactions, e.g., DKT uses LSTM to model complex student cognitive processes. Two variants of DKT have emerged in subsequent research. DKT+ Yeung and Yeung (2018) introduces two regularization terms to improve the consistency of KT predictions, while DKT-F Nagatani et al. (2019) enhances KT by considering forgetting behavior.
  • Attention-based models capture long-term dependencies between interactions through attention mechanisms, e.g., SAKT Pandey and Karypis (2019) is the first to use attention mechanisms to capture correlations between concepts and interactions. AKT Ghosh et al. (2020) employs a novel monotonic attention to represent the time distance between questions and students’ historical interactions. Due to AKT’s outstanding predictive performance, numerous powerful KT models are subsequently derived, such as simpleKT Liu et al. (2023b), FoLiBiKT Im et al. (2023), sparseKT Huang et al. (2023), and DTransformer Yin et al. (2023).
  • Mamba-based models are strong competitors to Transformer models. The recently proposed Mamba4KT Cao and Zhang (2024) is the first KT model to explore evaluation efficiency and resource utilization.
  • Graph-based models use graph structures to characterize the relationships between questions, concepts, or interactions, e.g., GKT Nakagawa et al. (2019) uses a graph to model the intrinsic relationships between concepts.
  • Memory-augmented models capture latent relationships between concepts through memory networks, e.g., DKVMN Zhang et al. (2017) uses a static key matrix to store relationships between different concepts and updates students’ knowledge states through a dynamic value matrix. SKVMN Abdelrahman and Wang (2019), a variant of DKVMN, also integrates the advantages of LSTM in recurrent modeling.
  • Adversarial-based models use adversarial techniques to enhance the model’s generalization ability, e.g., ATKT Guo et al. (2021) mitigates overfitting and improves generalization by adding perturbations to student interactions during training.
@@
  • Other representative models include interpretable models and models with auxiliary tasks, e.g., Deep-IRT Yeung (2019) introduces item response theory Lord (1952) based on DKVMN to make deep learning-based KT explainable. AT-DKT Liu et al. (2023a) enhances KT by introducing two auxiliary learning tasks: question tagging prediction and individualized prior knowledge prediction.
@@ 3.1 Problem Statement
3.2 Preliminary
LSTM [^2] Refer to Appendix A.2 for details on LSTM. is one of the earliest popular deep learning methods applied to NLP, but it has been overshadowed for a period by the success of Transformers Vaswani et al. (2017). However, the architecture is recently regaining attention and undergoing significant improvements. The improved LSTM is called extended Long Short-Term Memory (xLSTM) Beck et al. (2024), which mainly addresses three limitations in traditional LSTM: (1) inability to revise storage decisions, (2) limited storage capacities, and (3) lack of parallelizability. xLSTM introduces two new members to the LSTM family to overcome these limitations: sLSTM and mLSTM, as described in Figure 1. Since our work does not focus on the architecture of xLSTM, we have placed the detailed introduction of xLSTM in Appendix A.3.
@@ 3.3 DKT2
Figure  2 illustrates the architecture of our proposed DKT2, as described below.
Figure 2: Architecture of DKT2.
@@ 3.3.1 Rasch Embedding
We use the classic Rasch model Rasch (1993) from educational psychology to construct embeddings of questions and student skills. This model explicitly uses scalars to represent the degree of deviation between questions and the concepts they cover. Additionally, we choose to use question-specific difficulty vectors to capture differences among various questions within the same concept. DKT2 takes applicable interactions (i.e., inputs not involving the future response $r_{t+1}$, distinguishing it from models like AKT Ghosh et al. (2020)) as input, denoted as $\{q_{i},c_{i},r_{i}\}_{i=1}^{t}$, and at time step $t$, the embeddings of questions and student skills, $Q_{t}$ and $S_{t}$ respectively, are represented as:
$\begin{split}Q_{t}&=e_{c_{t}}+d_{q_{t}}\cdot\mu_{c_{t}},\\ S_{t}&=e_{(c_{t},r_{t})}+d_{q_{t}}\cdot g_{r_{t}},e_{(c_{t},r_{t})}=e_{c_{t}}+e_{r_{t}},\end{split}$ (2)
where $e_{c_{t}}\in\mathbb{R}^{d}$ and $e_{r_{t}}\in\mathbb{R}^{d}$ are the embeddings of concept $c_{t}$ and response $r_{t}$, respectively. $d_{q_{t}}\in\mathbb{R}$ is a difficulty scalar and $\mu_{c_{t}}\in\mathbb{R}^{d}$ summarizes the variation of questions containing concept $c_{t}$. $e_{(c_{t},r_{t})}\in\mathbb{R}^{d}$ is the interaction representation of the concept and student response (e.g., $c_{t}+r_{t}$), $g_{r_{t}}\in\mathbb{R}^{d}$ is the variant embedding of the response. $d$ is the dimension of the embeddings.
@@
[…]
$\left.\left.\left.\text{Res}\left(\text{LN}\left(\text{sLSTM}(S_{1:t})\right)\right)\right)\right)\right).$
where LN and Res refer to layer normalization Ba et al. (2016b) and residual connection He et al. (2015), respectively. Note that for simplicity, we have omitted operations like causal convolution from xLSTM. For detailed specifics, please refer to the settings described in the original paper (Beck et al., 2024).
@@ 3.3.3 IRT Prediction & Decompose
The core idea of IRT (Item Response Theory) lies in the interactive relationship between student ability and question difficulty Yen and Fitzpatrick (2006). Specifically, if a student’s ability is far above the question’s difficulty, the probability of the student responding to the question correctly is very high, and vice versa. This is also why IRT is often used for interpretable predictions in KT Yeung (2019); Sun et al. (2024). Therefore, the knowledge acquired by a student, denoted as $K_{1:t}$, can be represented as:
@@
$\begin{split}K_{1:t}^{+}&=\text{exp}(r_{1:t},d)\circ K_{1:t},\\ K_{1:t}^{-}&=\text{exp}(\textbf{one}-r_{1:t},d)\circ K_{1:t},\end{split}$ (5)

### v2

@@ Abstract
Knowledge Tracing (KT) is a fundamental component of Intelligent Tutoring Systems (ITS), enabling the modeling of students’ knowledge states to predict future performance. The introduction of Deep Knowledge Tracing (DKT), the first deep learning-based KT (DLKT) model, has brought significant advantages in terms of applicability and comprehensiveness. However, recent DLKT models, such as Attentive Knowledge Tracing (AKT), have often prioritized predictive performance at the expense of these benefits. While deep sequential models like DKT have shown potential, they face challenges related to parallel computing, storage decision modification, and limited storage capacity. To address these limitations, we propose DKT2, a novel KT model that leverages the recently developed xLSTM architecture. DKT2 enhances applicable input representation using the Rasch model and incorporates Item Response Theory (IRT) for output interpretability, allowing for the decomposition of learned knowledge into familiar and unfamiliar knowledge. By integrating this knowledge with predicted questions, DKT2 generates comprehensive knowledge states. Extensive experiments conducted across three large-scale datasets demonstrate that DKT2 consistently outperforms 18 baseline models in various prediction tasks, underscoring its potential for real-world educational applications. This work bridges the gap between theoretical advancements and practical implementation in KT. Our code and datasets are fully available at https://github.com/zyy-2001/DKT2.
Keywords:
Knowledge Tracing Information Interaction.
@@ 1 Introduction
The rapid expansion of educational data within Intelligent Tutoring Systems (ITS) [59] (e.g., AutoTutor [66]) has exposed significant limitations in traditional machine learning approaches [9]. In contrast, the advent of deep learning has introduced novel opportunities for addressing these challenges [45, 88]. A critical component of ITS is Knowledge Tracing (KT), which models students’ knowledge states and predicts future performance by analyzing their interaction data. Deep learning, with its advanced feature learning paradigm, offers enhanced modeling power and predictive accuracy in this context.
@@
Deep Knowledge Tracing (DKT) [69] represents the first significant application of deep learning to KT, employing Long Short-Term Memory (LSTM) networks [37] to capture the complexity of students’ learning processes. As a pioneering deep learning-based KT (DLKT) model, DKT has demonstrated superior predictive performance compared to traditional machine learning-based KT models (e.g., Bayesian Knowledge Tracing (BKT) [16]), offering notable advantages in applicability and comprehensiveness.
@@
DKT encodes students’ historical interactions to generate a comprehensive representation of their knowledge states (i.e., proficiency scores [^1] Proficiency scores range from 0 to 1, with higher values indicating greater knowledge and skill level. for each concept at each time step) and predicts future performance. However, recent DLKT models, such as the Attentive Knowledge Tracing (AKT) [26], while excelling in predictive accuracy [54, 42, 40, 95, 103], present limitations in applicability and comprehensiveness (the related details are in Sec. 3.2.2). Specifically, AKT requires both historical and future interactions as input, complicating its practical application since future responses are typically unavailable. Additionally, unlike DKT, AKT directly predicts scores on future questions without generating a comprehensive knowledge state, potentially weakening the correlations between different concepts and narrowing the definition of knowledge states in KT. Our review of 60 KT-model-related papers published in top AI/ML conferences and journals over the past decade (see Appendix 0.A.1) reveals a trend where evaluation performance has been prioritized at the expense of practical applicability, risking a disconnect between theoretical advancements and real-world implementation.
@@
Deep sequential models like DKT have intrinsic limitations that may prevent them from achieving optimal performance. LSTM networks, for instance, face challenges in dynamically updating stored information and exhibit limited storage capacity due to their scalar cell state design. Moreover, their inherent sequential processing nature hinders parallelization, limiting their scalability to large datasets. The recently proposed xLSTM [8], however, addresses these challenges by introducing two new variants: sLSTM, which improves LSTM’s storage decision by incorporating an exponential activation function, and mLSTM, which replaces scalar cell states with matrix memory for increased storage capacity and improved retrieval efficiency, while achieving full parallelization by abandoning memory mixing. Building on the strengths of xLSTM, we introduce DKT2, an enhanced DLKT model designed for greater applicability and comprehensiveness. DKT2 integrates the Rasch model [71] from educational psychology to process historical interactions, using xLSTM for knowledge learning. DKT2 then incorporates Item Response Theory (IRT) [58, 92] to interpret the learned knowledge, differentiating between familiar and unfamiliar knowledge, and ultimately integrates this knowledge with predicted questions to generate comprehensive knowledge states.
@@
  • We introduce DKT2, a model built on xLSTM, adhering to rigorous applicable input and comprehensive output settings, and incorporating both the Rasch model for input and an interpretable IRT-based output module.
  • We conduct extensive experiments, including one-step prediction, multi-step prediction, and predictions with varying history lengths, across three large-scale datasets. Our findings demonstrate that DKT2 consistently outperforms 18 baseline models, with additional analysis on the impact of input settings and multi-concept output predictions on KT performance.
@@ 2 Related Work
Since DKT [69] first applied deep learning methods to the KT task a decade ago, deep learning techniques have flourished in KT. Current DLKT models can be categorized into the following 8 types:
  • Deep sequential models use recurrent structures to encode students’ chronologically ordered interactions, e.g., DKT uses LSTM to model complex student cognitive processes. Two variants of DKT have emerged in subsequent research. DKT+ [94] introduces two regularization terms to improve the consistency of KT predictions, while DKT-F [64] enhances KT by considering forgetting behavior.
  • Attention-based models capture long-term dependencies between interactions through attention mechanisms, e.g., SAKT [67] is the first to use attention mechanisms to capture correlations between concepts and interactions. AKT [26] employs a novel monotonic attention to represent the time distance between questions and students’ historical interactions. Due to AKT’s outstanding predictive performance, numerous powerful KT models are subsequently derived, such as simpleKT [54], FoLiBiKT [42], sparseKT [40], DTransformer [95], and stableKT [47].
  • Mamba-based models are strong competitors to Transformer models. The recently proposed Mamba4KT [10] is the first KT model to explore evaluation efficiency and resource utilization.
  • Graph-based models use graph structures to characterize the relationships between questions, concepts, or interactions, e.g., GKT [65] uses a graph to model the intrinsic relationships between concepts.
  • Memory-augmented models capture latent relationships between concepts through memory networks, e.g., DKVMN [96] uses a static key matrix to store relationships between different concepts and updates students’ knowledge states through a dynamic value matrix. SKVMN [1], a variant of DKVMN, also integrates the advantages of LSTM in recurrent modeling.
  • Adversarial-based models use adversarial techniques to enhance the model’s generalization ability, e.g., ATKT [29] mitigates overfitting and improves generalization by adding perturbations to student interactions during training.
@@
  • Other representative models include interpretable models and models with auxiliary tasks, e.g., Deep-IRT [93] introduces item response theory [58] based on DKVMN to make deep learning-based KT explainable. AT-DKT [53] enhances KT by introducing two auxiliary learning tasks: question tagging prediction and individualized prior knowledge prediction.
@@ 3.1 Problem Statement
3.2 Preliminaries
3.2.1 LSTM and the Extended LSTM
LSTM [^2] Refer to Appendix 0.A.2 for details on LSTM. is one of the earliest popular deep learning methods applied to NLP, but it has been overshadowed for a period by the success of Transformers [82, 105]. However, the architecture is recently regaining attention and undergoing significant improvements. The improved LSTM is called extended Long Short-Term Memory (xLSTM) [8], which mainly addresses three limitations in traditional LSTM: (1) inability to revise storage decisions, (2) limited storage capacities, and (3) lack of parallelizability. xLSTM introduces two new members to the LSTM family to overcome these limitations: sLSTM and mLSTM, as described in Fig. 1. Since our work does not focus on the architecture of xLSTM, we have placed the detailed introduction of xLSTM in Appendix 0.A.3.
3.2.2 Weakly Applicable Input and Comprehensive Output Settings in DLKT Models
Figure 2: Structural sketch of AKT.
We use AKT [26] as an example to describe the common weakly applicable input and comprehensive output settings in DLKT models. Fig. 2 shows a structural sketch of AKT. Clearly, AKT takes both historical interactions and future interactions as input during training and inference, ignoring future information through attention masking while representing knowledge learned up to the current time step through offset (right-shifting values in attention), and directly predicts questions at each time step. From this, we can see that although AKT’s setup is reasonable and does not lead to future information leakage, this input setting, while convenient, also causes complications in engineering implementation (engineering often requires cumbersome representation of future information as a padding value, and this common processing method does not seem suitable for KT, as KT tasks typically involve predicting future questions $12\sim t+$ 1 based on historical interactions $1\sim t$). Moreover, AKT only outputs the response for the current time step’s question, without considering the student’s proficiency in different dimensions, which contradicts the multidimensional nature of real-world student knowledge and narrows the definition of KT.
Figure 3: Architecture of DKT2.
@@ 3.3 DKT2
Fig. 3 illustrates the architecture of our proposed DKT2, as described below.
@@ 3.3.1 Rasch Embedding
We use the classic Rasch model [71] from educational psychology to construct embeddings of questions and student skills. This model explicitly uses scalars to represent the degree of deviation between questions and the concepts they cover. Additionally, we choose to use question-specific difficulty vectors to capture differences among various questions within the same concept. DKT2 takes applicable interactions (i.e., inputs not involving the future response $r_{t+1}$, distinguishing it from models like AKT) as input, denoted as $\{q_{i},c_{i},r_{i}\}_{i=1}^{t}$, and at time step $t$, the embeddings of questions and student skills, $Q_{t}$ and $S_{t}$ respectively, are represented as:
$Q_{t}=e_{c_{t}}+d_{q_{t}}\cdot\mu_{c_{t}},S_{t}=e_{(c_{t},r_{t})}+d_{q_{t}}\cdot g_{r_{t}},e_{(c_{t},r_{t})}=e_{c_{t}}+e_{r_{t}},$ (2)
where $e_{c_{t}}\in\mathbb{R}^{d}$ and $e_{r_{t}}\in\mathbb{R}^{d}$ are the embeddings of concept $c_{t}$ and response $r_{t}$, respectively. $d_{q_{t}}\in\mathbb{R}$ is a difficulty scalar and $\mu_{c_{t}}\in\mathbb{R}^{d}$ summarizes the variation of questions containing concept $c_{t}$. $e_{(c_{t},r_{t})}\in\mathbb{R}^{d}$ is the interaction representation of the concept and student response, $g_{r_{t}}\in\mathbb{R}^{d}$ is the variant embedding of the response. $d$ is the dimension of the embeddings.
@@
$A_{1:t}=\text{Res}\left(\text{LN}\left(\text{mLSTM}\left(\text{Res}\left(\text{LN}\left(\text{sLSTM}(S_{1:t})\right)\right)\right)\right)\right),$ (3)
where LN and Res refer to layer normalization [7] and residual connection [31], respectively.
@@ 3.3.3 IRT Prediction & Decompose
The core idea of IRT (Item Response Theory) lies in the interactive relationship between student ability and question difficulty [92]. Specifically, if a student’s ability is far above the question’s difficulty, the probability of the student responding to the question correctly is very high, and vice versa. This is also why IRT is often used for interpretable predictions in KT [93, 79] (our work focuses not on the interpretability of KT models but on evaluating their applicability and comprehensive setup). Therefore, the knowledge acquired by a student, denoted as $K_{1:t}$, can be represented as:
@@
$\begin{split}K_{1:t}^{+}=\text{exp}(r_{1:t},d)\circ K_{1:t},K_{1:t}^{-}=\text{exp}(\textbf{one}-r_{1:t},d)\circ K_{1:t},\end{split}$ (5)
@@
3.3.4 Integrated Knowledge Fusion
Finally, DKT2 predicts the student’s comprehensive knowledge states $\text{KS}_{2:t+1}$:
$\begin{split}\text{KS}_{2:t+1}=\sigma(\text{ReLU}(X_{2:t+1}W_{1}+b_{1})W_{2}+b_{2}),\end{split}$ (7)
Eq. 7 can be further represented as:
$\begin{split}\text{KS}_{i}=(\hat{r}_{i}^{1},\hat{r}_{i}^{2},\dots,\hat{r}_{i}^{n}),2\leq i\leq t+1,\end{split}$ (8)
where $\hat{r}_{i}^{j}$ represents the prediction score of DKT2 for concept $j$ at time step $i$. The comprehensive output of DKT2 enables the prediction of multiple concepts at the same time step, whereas models like AKT can only predict $\hat{r}_{i}^{c_{i}}$ at time step $i$. We will analyze the multi-concept prediction scenario in detail in Sec. 4.3.2, where some unexpected results have been discovered.
3.3.5 Model Training
$\mathcal{L}_{\text{DKT2}}=-\sum_{i=2}^{t+1}{r_{i}\text{log}({\hat{r}}_{i})+(1-r_{i})\text{log}(1-{\hat{r}}_{i})}.$ (9)
3.3.6 Conversion of Input and Output Settings
We attempt to convert the weakly applicable input and comprehensive output settings in DLKT models into strongly applicable input and comprehensive output settings. Similarly, using AKT as an example, like DKT2, as shown in Fig. 3, the transformed AKT only takes the historical interactions $\{(q_{i},c_{i},r_{i})\}_{i=1}^{t}$ as input, with everything else remaining unchanged (note that the right-shift operation still needs to be retained because the attention does not mask the knowledge of the current time step). Before outputting the predicted score $\hat{r}_{i}^{c_{i}}$, it first concretizes knowledge into the knowledge of each concept (by converting the original dimensions into the number of concepts through an MLP) and then the comprehensive knowledge state is obtained through a Sigmoid function.
4 Experiments
Our goal is to answer the following research questions:
  • RQ1: How does DKT2 perform compared to 18 baselines from 8 different categories under applicable input and comprehensive output settings?
4.1 Experimental Setup
4.1.1 Datasets
We conduct extensive experiments on three of the latest large-scale benchmark datasets from different platforms: Assist17 [23], EdNet [15], and Comp [38]. Details of the datasets are provided in Appendix 0.A.4.
4.1.2 Baselines
To comprehensively and systematically evaluate the performance of DKT2 and analyze the impact of input-output settings on KT models, we compare DKT2 with 18 DLKT baselines from 8 categories, as mentioned in Sec. 2. Detailed descriptions of the aforementioned DLKT baselines can be found in Appendix 0.A.5.
4.1.3 Implementation
Similar to CL4KT [46], we employ five-fold cross-validation, with folds divided by students. 10% of the training set is used for model evaluation and also for the early stopping strategy: if the AUC does not improve within 10 epochs during the 300 epochs, the training will be stopped. The averages across five test folds are reported. We focus on the most recent 100 interactions (history length) for each student, as this latest information is crucial for future predictions. During training, all models are trained using the Adam optimizer [43] with the following settings: batch size is fixed at 512, learning rate is 0.001, dropout rate is 0.05, and embedding dimension is 64. The seed is set to 12405 to reproduce experimental results. Similar to existing DLKT research, our evaluation metrics include two classification metrics, Area Under the ROC Curve (AUC) and Accuracy (ACC), and one regression metric, Root Mean Square Error (RMSE). Note that our experimental parameter configuration is consistent with CL4KT.
4.2 Applicable and Comprehensive Performance Comparison (RQ1)
Under applicable input and comprehensive output settings, we evaluate three common prediction tasks in KT [55]: 1) one-step prediction, 2) multi-step prediction, and 3) prediction with varying history lengths.